# “You Might Like It”: How People Respond to Small Talk in Human-Robot Collaboration

### v1

@@ 5.4. Limitations and Future Work
Despite its insights, our study’s short duration and single-session focus may not fully capture the depth, intricacies and dynamics of extended human-robot collaborations and how they are influenced by small talk. Additionally, the study did not consider cultural nuances of participants, which are vital as different cultures perceive and value small talk differently. While this work serves as an exploratory examination of the use of small talk in human-robot interactions, future studies should focus on testing hypotheses to quantify the benefits of small talk for rapport building and to examine its role in human-robot collaborations with negative feedback, team evaluations, or robot errors. When designing a conversational framework for robot small talk, it is essential to research and ensure the appropriateness of content, timing, and context. For instance, the framework should be sensitive to user preferences on topics, adapt according to the user’s state for ideal timing, and consider team performance metrics to prevent disruptions during critical tasks. Lastly, future research should explore methods to align a robot’s conversational capabilities with its functional abilities. It is important to keep users apprised of the robot’s true abilities, possibly using small talk as an informative medium. Nevertheless, the strategic introduction of small talk can be instrumental in building and strengthening rapport between users and task-oriented robots in collaborative scenarios.
@@ Acknowledgements.
Table 3. Overview of Codes and Their Definitions for the Data Analysis
@@
Answer
Response (Yes) and answered robot’s question
Refuse
Response (Yes) but did not answer the robot’s question
No engagement
Participant did not answer the robot’s question
@@
Minimal (glance)
Looks at the robot but no task disruption
Freeze
Stops current action (minimum of 2 seconds) and looks at the robot before continuing the task
None
Did not stop or look at the robot
@@
Pleasant
Individual exhibits a smile
Surprise
Individual exhibits behaviors such as a chin raise, lower lip raise, brow lowering, etc.
Indifference
No reaction
Non-Verbal Reactions to Negative Feedback
Name Definition
Verbal Response
An audible or verbal phrase spoken by the participant
Non-Verbal Response
Behaviors such as smiling, laughing, eye-rolls, head nods, arm movements, etc.
Blame Attribution
Name Definition
Self
Selects themselves as a potential cause for the negative feedback announcement
Robot
Selects the robot as a potential cause for the negative feedback announcement
Team
Selects the team as a potential cause for the negative feedback announcement
Nobody
Believes neither they nor the robot caused the negative feedback announcement
Motives for Initiating Questions
Name Definition
Curiosity
Individual describes a level of curiosity
Felt Human-like
Individual makes connections to human-likeness
Social Convention
Individual describes factors relating to social conventions
l6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 5.3. Limitations and Future Work
Despite its insights, our study’s short duration and single-session focus may not fully capture the depth, intricacies, and dynamics of extended human-robot collaborations and how they are influenced by small talk. This work used a small sample size $(N=20)$ and only studied one-time interactions. Thus, prolonged effects are unclear, and future work is needed to investigate longer or multi-session interactions. Additionally, the study did not consider any correlations between participants’ cultural nuances and their small talk engagements; this is vital as separate cultures perceive and value small talk differently (Cui, 2015; Rings, 1994; Bickmore and Cassell, 2005). The study also did not accurately evaluate its intended target demographic, as participants were recruited from a local university campus and surrounding areas instead of focusing on industrial or warehouse workers. Moreover, it was conducted in a lab setting with a low-stakes, simulated warehouse task. While this work examines the use of small talk in human-robot collaboration, future studies should focus on quantitatively evaluating the benefits of small talk for rapport building, how the user’s personality, cultural background, and task context may affect their small talk interaction, and the role of small talk with cognitively demanding user tasks. Ensuring the appropriateness of content, frequency, timing, and context are essential in designing conversational aspects of robot small talk in future research. For instance, a robot conversational system could be sensitive to user preferences on topics, adapt to individual users for ideal timing, and consider performance metrics to prevent disruptions during critical tasks; it is necessary to study the effects of different aspects of a robot’s conversational characteristics. Lastly, future research should explore methods to align a robot’s conversational capabilities with its functional abilities. It is important to keep users apprised of the robot’s true abilities, possibly using small talk as an informative medium.
6. Conclusion
This study explored user engagement with robots that initiate small talk in a simulated industrial collaborative setting. Our findings revealed a general willingness among users to engage in small talk with robots. Although users are receptive to robotic systems with small talk capabilities, which suggests potential enhancements to interaction experiences, we lack concrete evidence of the measurable benefits of robot-initiated small talk. Important interaction characteristics were identified, emphasizing the need for flexibility in the robot systems to adapt small talk’s frequency, complexity, and content to suit various users and tasks. However, introducing small talk could potentially distract users and reduce task focus and productivity, particularly during critical tasks. Therefore, limiting or avoiding small talk in such situations might be prudent. While social chat with robots is not new, integrating robot small talk in physical, collaborative settings with humans is a newer concept. Given the increasing prevalence of such robots in real-world settings and the existing interest in social interaction among industrial workers, we recommend further research to implement and quantify the benefits of social conversations between humans and robots in these environments.
@@ Acknowledgements.
Table 6. Overview of Codes and Their Definitions for the Data Analysis: Part A
Robot’s Verbal Phrases: Task-Oriented
Name Definition
Confirmation Robot confirms it processed user request
Reminder Robot gives a reminder message about the task
Robot’s Verbal Phrases: Small-Talk
Name Definition
Question Robot initiates a question to the user
Statement Robot makes a small-talk response or comment
Participant’s Verbal Phrases: Task-Oriented
Name Definition
Request User requests a yellow or green pipe, or reports a faulty joint
Reminder User responses to the robot’s reminder message
Confirmation User responses to the robot’s confirmation message
System Announcement User responses to the negative system announcement message
Participant’s Verbal Phrases: Small Talk
Name Definition
Question User initiates a question to the robot
Repeat User asks the robot to repeat or clarify what it just said
Short Non-question small talk statement ¡= 5 words
Medium Non-question small talk statement ¿ 5 words
Long Multi-sentence non-question small talk statements
Table 7. Overview of Codes and Their Definitions for the Data Analysis: Part B
Motives for Initiating Questions
Name Definition
Curiosity Individual describes a level of curiosity
Felt Human-like Individual makes connections to human-likeness
Social Convention Individual describes factors relating to social conventions
Non-Verbal Reactions to Negative Feedback
Name Definition
Verbal Response An audible or verbal phrase spoken by the participant
Non-Verbal Response Behaviors such as smiling, laughing, eye-rolls, head nods, etc.
Blame Attribution
Name Definition
Self Selects themselves as a potential cause for the feedback announcement
Robot Selects the robot as a potential cause for the feedback announcement
Team Selects the team as a potential cause for the feedback announcement
Nobody Believes neither they nor the robot caused the feedback announcement
@@
Answer Response (Yes) and answered robot’s question
Refuse Response (Yes) but did not answer the robot’s question
No engagement Participant did not answer the robot’s question
@@
Minimal (glance) Looks at the robot but no task disruption
Freeze Stops current action (min. 2 sec) and looks at robot before continuing
None Did not stop or look at the robot
@@
Pleasant Individual exhibits a smile
Surprise Individual exhibits behaviors such as a chin raise, lower lip raise, brow lowering, etc.
Indifference No reaction